Square root 8-x -x=4
Answer by CubeyThePenguin(3113)

sqrt(8 - x) = 4 + x
8 - x = x^2 + 8x + 16
0 = x^2 + 9x + 8
0 = (x + 1)(x + 8)

x = -1 or x = -8 (extraneous)

x+=+-1